The Eastern Division of Acme, Inc. has operating income of $32,000 on sales revenue of $320,000. Divisional operating assets are $160,000, and management of Acme has determined that a minimum return of 10% should be expected from all investments.

a. Using the DuPont model, calculate Acme's margin, turnover, and ROI.

b. Calculate Acme's residual income

c. Should Acme accept a project that is expected to achieve a 15% ROI if it is being evaluated using residual income?
Answer #1

Margin = Net operating income/sales

= 32,000/320,000

= 10%

Turnover = Sales/Operating assets

= 320,000/160,000

= 2

ROI = margin * turnover

= 10*2

= 20%

Residual income = Operating income - Required return

= 32,000 - (160,000*10%)

= 16,000

C.